OpenStudy (anonymous):

1/y-2=2/y+1 what is y?

OpenStudy (amistre64):

what have you tried and come up with?

OpenStudy (anonymous):

i tried -1/3 and that didnt work

OpenStudy (amistre64):

i take it those are y-2 and y+1 in the denominators?

OpenStudy (anonymous):

yes they are

OpenStudy (amistre64):

we can start this by something they call "cross multiplying" which is more descriptive than mathical

OpenStudy (amistre64):

take the bottom of one side and multiply it to the top of the other side

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

1 2 ---- = ---- --> 1(y+1) = 2(y-2) y-2 y+1

OpenStudy (amistre64):

what should our next step be?

OpenStudy (amistre64):

Hannah: now that we have this 1(y+1) = 2(y-2) what do you think we can do with it?

OpenStudy (anonymous):

combine them?

OpenStudy (amistre64):

combine is a vague term, but its good that you are participating :) lets multiply those numbers that are outside the (..)s thru like this: 1(y+1) = 2(y-2) y+1 = 2y -4 now that its all adding; how do we get y all alone?

OpenStudy (anonymous):

subtract y?

OpenStudy (amistre64):

very good :) that is a step in the right direction for this y+1 = 2y -4 -y -y ------------ 1 = y - 4 ; now how should we get rid of the -4?

OpenStudy (anonymous):

add it!

OpenStudy (amistre64):

yes!! :) 1 = y -4 +4 +4 --------- 5 = y ; and thats our answer. Good job!

OpenStudy (anonymous):

thank you!!

OpenStudy (amistre64):

youre welcome :)